if you randomly select a card from a standard deck (52 cards) find the odds against it being a black face card
1
Expert's answer
2012-03-13T10:34:12-0400
so, let event A - selected card is black
There are 26 black faced cards in the standard deck

So, P(A)=26/52=1/2